while standing in a low tunnel you push up on the ceiling with a force of 100N. your mass is 70kg. What force does the floor exert on you?
Answer #1
A) I'm with you, the force from the ceiling is 100 N.

B) The floor supports both the 100N force applied to the ceiling, and the weight of the person.

F = m*g + 100 N =70*9.8 + 100 = 786 N, rounded to 790 N.
Answer #2
100 + 70*9.8= 756 N